Business and finance. Kevin earned $165 interest for 1 year on an investment of
$1500. At the same rate, what amount of interest would be earned by an investment of $2500?

Answer by checkley75(3666)

165/1500=X/2500 CROSS MULTIPLYING WE GET:
1500X=165*2500
1500X=412500
X=412500/1500
X=$275 IS THE INTEREST ON AN INVESTMENT OF $2500.
BASED ON AN INTEREST RATE OF 11% (165/1500=.11=11%).